caira12:

Solve for x: 1/8 (8x + 15) = 24 (5 points) Select one: a. 9 b. 49 c. 177/8 d. 207/8

Vocaloid:

notice how the whole (8x+15) is being multiplied by (1/8) any ideas how we would "undo" this and "eliminate" the 1/8?

caira12:

ummm. nr

Vocaloid:

so to "undo" it we would do the opposite of multiplying by (1/8) so you would divide each side of the equation by 1/8

Vocaloid:

can you try taking the equation 1/8 (8x + 15) = 24 and dividing both sides by (1/8)?

caira12:

i will try

caira12:

wait, wdym both sides?

Vocaloid:

when you have an equation like A = B the left side is A and the right side is B

caira12:

oh

Vocaloid:

so if we asked "divide both sides by (1/8)" we would get A/(1/8) = B/(1/8)

Vocaloid:

try to apply this same logic to 1/8 (8x + 15) = 24

caira12:

so

Vocaloid:

(1/8)(8x+15)/(1/8) = ? as a hint, the 1/8 in the numerator and the 1/8 in the denominator cancel out

caira12:

the final answer would be A

Vocaloid:

hm not quite

Vocaloid:

(1/8)(8x+15)/(1/8) = ?

caira12:

ohhhhhh, 1/8 = 8

Vocaloid:



Vocaloid:



Vocaloid:

notice how the 1/8 in the numerator and denominator cancel out so (1/8)(8x+15)/(1/8) = 8x + 15 now apply this same logic to the other side of the equation 24/(1/8) = ?

caira12:

24(1/8)=3

Vocaloid:

be careful, it's 24 divided by (1/8) not 24 * 1/8

Vocaloid:

if it helps, 24/(1/8) is the same thing as 24*8

caira12:

ohhhhh so it would be 192

Vocaloid:

good so 8x + 15 = 192 since we have 15 added to 8x, we can do the opposite of adding 15 to isolate 8x , what's the opposite of adding 15?

caira12:

subtracting 15

Vocaloid:

good so take 8x + 15 = 192 and subtract 15 from both sides

jhonyy9:

courage @caira12 this is easy now

caira12:

final answer is 177/8

Vocaloid:

good